# FLUXSynID: A Framework for Identity-Controlled Synthetic Face Generation with Document and Live Images

### v2

@@ Abstract
Synthetic face datasets are increasingly used to overcome the limitations of real-world biometric data, including privacy concerns, demographic imbalance, and high collection costs. However, many existing methods lack fine-grained control over identity attributes and fail to produce paired, identity-consistent images under structured capture conditions. We introduce FLUXSynID, a framework for generating high-resolution synthetic face datasets with user-defined identity attribute distributions and paired document-style and trusted live capture images. The dataset generated using the FLUXSynID framework shows improved alignment with real-world identity distributions and greater inter-set diversity compared to prior work. The FLUXSynID framework for generating custom datasets, along with a dataset of 14,889 synthetic identities, is publicly released [^1] https://github.com/Raul2718/FLUXSynID to support biometric research, including face recognition and morphing attack detection.
@@ 1 Introduction
Large-scale, diverse, and well-annotated face datasets are crucial for training and evaluating biometric systems. In addition to face recognition [40, 44, 49] and emotion analysis [51, 15], recent work on morphing attack detection (MAD) [7, 34, 2] and face demorphing [35, 53] also depends heavily on such datasets. However, collecting real-world facial data presents substantial challenges. Privacy regulations restrict access and usage, while the high financial cost of acquiring and annotating large volumes of data can be prohibitive. Moreover, demographic imbalances and the need for operationally relevant conditions, such as capturing both document-style and trusted live images for the same subject, further complicate data collection efforts.
@@
While generative models such as Variational Autoencoders (VAEs) [30], GANs [20], and Diffusion Models [50] have enabled synthetic face dataset creation [39, 5, 45, 6, 27], most existing methods focus on in-the-wild imagery and lack support for structured capture conditions or fine-grained control over identity attributes (e.g., age, ethnicity). Moreover, generated faces often diverge from the distributional characteristics of real data [25]. The ONOT dataset [14] addresses some of these issues but remains limited in scale, control, and intra-set diversity.
@@
In this work, we introduce FLUXSynID, a framework for generating synthetic face datasets with controllable identity attributes and paired document-style and live capture images. Built on the FLUX.1 [dev] diffusion model [3] and enhanced via LoRA-based fine-tuning [24], FLUXSynID produces high-quality, identity-consistent images guided by natural language prompts. It employs a two-stage pipeline: generating document-style images first, then corresponding live capture variants using three complementary methods. These are based on LivePortrait [21], PuLID [22], and Arc2Face [42], each introducing different levels of variation in pose, expression, lighting, and background.
The contributions of this work are threefold:
@@
  • A method for generating high-quality document-style face images with user-defined identity attribute distributions, allowing direct demographic control.
  • A multi-method approach to generating synthetic trusted live capture images that preserve identity while introducing realistic capture variability.
@@
  • A comparative evaluation of FLUXSynID against comparable real (FRGC [43], FRLL [11]) and synthetic (ONOT [14], StyleGAN2 [27]) face datasets, demonstrating improved alignment in visual appearance and in embedding space distributions produced by face recognition systems.
@@
To support research in morphing attack detection and related biometric fields, we publicly release the FLUXSynID framework along with a dataset of 14,889 synthetic identities.
@@ 2 Related Work
Synthetic face datasets are increasingly used in biometric research due to the limitations of real facial data, such as privacy issues, high costs, and demographic imbalances [25, 39, 14, 5]. Early methods such as SynFace [45] and DigiFace-1M [1] showed promise but lacked realism and flexibility. SynFace used a 3D-aware GAN to generate facial images with limited variation, while DigiFace-1M used 3D rendering to create large-scale identities but struggled to match real-world distributions and offered limited control over attributes such as age. Both fell short in visual fidelity and attribute precision, limiting their use in tasks such as morphing attack detection (MAD).
GANDiffFace [39] is a recent approach combining two generative models to achieve explicit control over demographics and identity. It uses StyleGAN3 [26] to generate high-resolution synthetic faces and applies linear Support Vector Machines (SVMs) for targeted manipulation of demographic attributes. To introduce intra-class variation, a diffusion model is fine-tuned using DreamBooth [52] on each generated identity, allowing for controlled sampling. Despite these advancements, GANDiffFace, similar to SFace [5], focuses on unconstrained images and lacks document-style outputs. StyleGAN3, while powerful, struggles with consistent identity preservation across variations and requires a separate SVM for each attribute. Furthermore, the need to fine-tune a diffusion model per identity adds significant computational overhead, limiting the scalability of this approach for biometric tasks.
Recent studies have leveraged diffusion models for identity-consistent face generation, achieving greater visual realism and intra-class variation. DCFace [29] introduced a dual-conditioning approach using both identity and style embeddings to generate varied images of the same person. IDiff-Face [4] extended this by conditioning a latent diffusion model on face embeddings from recognition networks, producing distinct yet consistent identities. Both demonstrate strong recognition performance and show that diffusion models better capture real-world identity spaces. However, they focus on unconstrained synthesis and do not support document-style imagery paired with live captures. Additionally, identity attribute control is limited to what the embeddings encode, lacking mechanisms for explicit demographic customization.
The ONOT dataset [14] is the only prior work focused on document-style face synthesis for biometric use. It employed a fine-tuned Stable Diffusion [50] model with prompt engineering to generate ICAO-compliant [56] identity images, sampling 64 images per identity across 15k identities and retaining only those with at least one compliant image. However, this brute-force filtering resulted in usable samples for just 27% of identities, with further intra- and inter-class filtering reducing the dataset to a maximum of 255 identities. Our analysis shows that even these retained identities were highly similar and poorly aligned with the real identity embedding space. ONOT’s document-style control relies solely on prompts, lacking the fine-tuning necessary for consistent ID-photo outputs. Its demographic attribute control is limited and not easily extensible to new or composite features. While ONOT shares our high-level goals, it falls short in flexibility, scalability, and representational diversity.
@@ 3 FLUXSynID: Document Images
Figure 1: Overview of the FLUXSynID pipeline for synthetic document image generation. Given user-defined identity attributes and constraints, descriptive prompts are generated and used to guide FLUX.1 [dev] [3] diffusion model enhanced with LoRA adapter [24] for producing realistic frontal ID-like images.
Figure 1 presents an overview of the framework for the generation of synthetic document images in FLUXSynID dataset. The FLUX.1 [dev] [3] diffusion model is used to generate high-quality, high-resolution images based on user-defined identity attributes.
@@ 3.1 Identity Attributes
Depending on the application, existing face image datasets may exhibit imbalances in identity attributes such as gender, age, or ethnicity [25]. To address this, the FLUXSynID framework allows user-defined attribute classes, each with a set of attributes with associated selection probabilities. Additionally, each class can be assigned an inclusion probability during sample generation. As shown in Figure 1, the gender class is always included (100%), with the Female identity attribute selected 49% of the time. Other classes, such as body type, may be included in only a portion of samples (e.g., 50%), enabling fine-grained control over attribute distribution in the synthetic dataset.
However, some attribute combinations can lead to semantic inconsistencies. For example, assigning Bald hairstyle along with a hair color such as Black creates a contradiction. To address this, FLUXSynID allows the definition of attribute clashes. Users can specify primary and secondary attributes, where the primary is retained and conflicting secondaries are discarded. In the above example, if Bald hairstyle is selected, any hair color attribute is excluded to ensure consistency.
Once a set of attributes is selected for a synthetic identity, it is stored for future reference, and a Large Language Model (LLM) then generates a coherent textual prompt for image synthesis. The use of an LLM is motivated by the fact that FLUX.1 [dev] [3] employs a T5 [47] encoder, which excels at capturing nuanced semantic relationships from verbose, detailed prompts. Specifically, the pre-trained Qwen2.5 [57] model is employed to generate natural language descriptions that consistently and semantically integrate the selected attributes.
@@ 3.2 Document Image LoRA Adapter
Low-Rank Adaptation (LoRA) [24] enables efficient fine-tuning of large pre-trained models by introducing a small set of trainable parameters. Instead of updating the full model weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, the LoRA adapter keeps $W_{0}$ fixed and learns a low-rank update $\Delta W=BA$, where $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times k}$, with $r\ll\min(d,k)$. The adapted weights become $W=W_{0}+\Delta W$. This approach significantly reduces trainable parameters compared to methods such as DreamBooth [52], used in generating the GANDiffFace [39] dataset.
In this work, a LoRA adapter is used to fine-tune the FLUX.1 [dev] [3] diffusion model for stylistic consistency in image generation. The goal is to generate high-resolution, vivid frontal face images with neutral expressions and a consistent white background. The adapter is trained on images from the Chicago Face Database (CFD) [37, 38, 32], which consists of high-quality portraits that match these criteria. Figure 2 shows outputs with and without LoRA, highlighting improved pose, lighting, and expression consistency.
Figure 2: Comparison of synthetic document images generated with and without LoRA [24]. Without LoRA, images exhibit inconsistent head poses, facial expressions, and lighting conditions. LoRA fine-tuning produces more uniform, ID-style images with neutral expressions and consistent photographic conditions.
@@ 3.3 Image Generation
Once trained, the LoRA adapter is used to update the weights of the pre-trained and frozen FLUX.1 [dev] [3] model, which conditions image generation based on prompts produced by Qwen2.5 [57]. FLUX.1 [dev] is a rectifying flow transformer [16] latent diffusion [50] model tailored for text-to-image synthesis. As a guidance-distill model, FLUX.1 [dev] incorporates Classifier-Free Guidance (CFG) [23] directly into the transformer through a modulation scale referred to as a guidance scale, eliminating the need for negative prompts and halving inference steps by avoiding separate conditional and unconditional passes. The guidance scale directly controls prompt adherence: lower values result in more diverse outputs with weaker attribute fidelity, while higher values tend to produce oversaturated, less varied images.
As shown in Figure 1, the generated prompt $P$ is encoded using the pre-trained and frozen CLIP-L [46] and T5-XXL [47] models. FLUX.1 [dev] leverages these encoders complementarily: CLIP’s global embedding is combined with the timestep embedding and used to modulate transformer layers, while T5-XXL’s token embeddings are concatenated with latent image tokens to form a unified sequence processed via self-attention. This achieves effects similar to cross-attention, allowing text tokens to guide specific image regions [19].
After encoding the text prompt, $P$, using both encoders ($E_{\text{CLIP-L \& T5-XXL}}$), the resulting textual conditioning, $c_{P}$, is used as conditional input to FLUX.1 [dev] alongside a 16-channel Gaussian noise latent image $z_{t}$ sampled from a random seed. Through latent diffusion, FLUX.1 [dev] iteratively denoises $z_{t}$ over $t$ steps to produce the final latent image. The conditional denoising process is defined as:
@@
where DPM++ 2M [36] with the beta schedule [33] is used as the sampling algorithm in diffusion probabilistic modeling, and $\epsilon_{\theta^{CFD}}$ represents the FLUX.1 [dev] model with the LoRA adapter trained on CFD images (see Section 3.2).
The final denoised latent image, $z_{0}$ is then passed to the pre-trained FLUX.1 [dev] Variational Autoencoder (VAE) decoder network [3], $D_{\text{VAE}}$, which generates the synthetic document image, $I_{D}$, containing the identity attributes described in the textual prompt.
@@ 4 FLUXSynID: Trusted Live Capture Images
Figure 3: Overview of the FLUXSynID live capture image generation pipeline. Starting from a synthetic document image, three methods are used to produce identity-consistent live images: LivePortrait [21] enables controlled variations in pose and expression; Arc2Face [42] introduces natural diversity via identity embeddings; and the PuLID-based approach [22, 54] is integrated with FLUX.1 [dev] [3] via identity and edge-based conditioning, guided by modified image prompts.
Figure 3 shows the framework for generating synthetic trusted live capture images for the FLUXSynID dataset. Three methods are used: LivePortrait [21], Arc2Face [42], and PuLID [22], producing live images $I_{LL}$, $I_{LA}$, and $I_{LP}$, respectively. Each method provides unique strengths in identity preservation, expression, lighting, and pose variation, making them well-suited for diverse applications.
@@ 4.1 LivePortrait-based Image Generation
LivePortrait [21] is a facial animation method that modifies facial expressions and head orientations from a single input image. Unlike diffusion-based methods, it employs an implicit-keypoint approach that allows for greater efficiency and fine-grained control. Its retargeting mechanism employs compact keypoints to enable subtle adjustments to facial features and pose, with a stitching module that ensures visual coherence by seamlessly blending altered facial regions.
Within FLUXSynID, LivePortrait generates modified versions of document images, introducing controlled variations randomly sampled for each image. This preserves identity while subtly adjusting pose and expression.
@@ 4.2 Arc2Face-based Image Generation
Arc2Face [42] is an identity-conditioned image generation framework that synthesizes realistic facial images using identity embeddings extracted via ArcFace [12]. It adapts a pre-trained Stable Diffusion [50] model to generate images directly from identity vectors, ensuring strong identity preservation. Unlike LivePortrait [21], Arc2Face introduces natural variations in pose, expression, and lighting without explicit user control. The authors report superior identity similarity compared to other generative models, underscoring its effectiveness. Notably, Arc2Face requires no textual prompts, relying entirely on image-based identity embeddings.
In FLUXSynID, Arc2Face is used to generate synthetic images directly from ArcFace embeddings of the original synthetic document images. This introduces natural diversity in expressions, poses, and lighting while maintaining identity consistency. The method complements the more controlled LivePortrait approach, enriching the dataset with greater variability.

### v3

@@ Abstract
Synthetic face datasets are increasingly used to overcome the limitations of real-world biometric data, including privacy concerns, demographic imbalance, and high collection costs. However, many existing methods lack fine-grained control over identity attributes and fail to produce paired, identity-consistent images under structured capture conditions. We introduce FLUXSynID, a framework for generating high-resolution synthetic face datasets along with a dataset of 14,889 synthetic identities. We generate synthetic faces with user-defined identity attribute distributions, offering both document-style and trusted live capture images. The dataset generated using the FLUXSynID framework shows improved alignment with real-world identity distributions and greater inter-class diversity compared to prior work. Our work is publicly released [^1] https://github.com/Raul2718/FLUXSynID to support biometric research, including face recognition and morphing attack detection.
@@ 1 Introduction
Large-scale, diverse, and well-annotated face datasets are crucial for training and evaluating biometric systems. In addition to face recognition [38, 42, 47] and emotion analysis [49, 15], recent work on Morphing Attack Detection (MAD) [7, 32, 2] and face demorphing [33, 51] relies significantly on such datasets. However, collecting real-world facial data presents substantial challenges. Privacy regulations restrict access and usage, while the high financial cost of acquiring and annotating large volumes of data can be prohibitive. Moreover, demographic imbalances and the need for operationally relevant conditions, e.g., capturing both document-style and trusted live capture images for the same subject, further complicate data collection efforts.
In particular, applications such as Differential Image-Based MAD (D-MAD) [54] require datasets containing paired, identity-consistent document-style and live capture images. In typical scenarios, such as border control, the document image may be a bona fide or morphed photo, while the live capture serves as a trusted reference. These tasks also require demographically diverse identities to ensure fairness and robustness. Yet, existing datasets are often limited in scale, and collecting real data with such properties raises legal, ethical, and logistical challenges.
@@
While recent advances in generative models have facilitated the creation of synthetic datasets [37, 4, 43, 5, 26], most focus on in-the-wild imagery and lack support for structured capture conditions or fine-grained control over identity attributes (e.g., age, ethnicity). Moreover, generated faces often diverge from the distributional characteristics of real data [24]. The ONOT dataset [14] addresses some of these issues but remains limited in scale, control, and inter-class diversity.
@@
In this work, we introduce FLUXSynID, a framework for generating synthetic face datasets with controllable identity attributes and paired document-style and live capture images. Built on the FLUX.1 [dev] [3] diffusion model and enhanced via LoRA-based fine-tuning [23], FLUXSynID produces high-quality, identity-consistent images guided by natural language prompts. It employs a two-stage pipeline: generating document-style images first, then corresponding live capture variants using three complementary methods. These are based on LivePortrait [20], Arc2Face [40], and PuLID [21], each introducing different levels of variation in pose, expression, lighting, and background.
The contributions of this work are as follows:
@@
  • A method for generating high-quality document-style face images with user-defined identity attribute distributions, enabling direct demographic control.
  • A novel live image generation approach that adapts the FLUX.1 [dev] and PuLID frameworks with new conditioning strategies for identity and structure preservation. Two complementary methods, LivePortrait and Arc2Face, are also integrated to generate multiple, diverse face images per identity, supporting downstream tasks such as D-MAD.
@@
  • A comparative evaluation against real and synthetic face datasets, demonstrating closer alignment of FLUXSynID with real data in both visual appearance and embedding-space distributions. The demographic bias of the generative model is also investigated and reported.
@@
  • The public release of the FLUXSynID framework and a dataset of 14,889 synthetic identities to support biometric research.
@@ 2 Related Work
Synthetic face datasets are increasingly used in biometric research due to the limitations of real facial data, such as privacy concerns, high costs, and demographic imbalances [24, 37, 14, 4]. SynFace [43] and DigiFace-1M [1] showed promise but offered limited variation, realism, and identity control, restricting their applicability to tasks such as Morphing Attack Detection (MAD) [54].
GANDiffFace [37] combines StyleGAN3 [27] with linear SVMs to control demographic attributes in synthetic face generation, and introduces intra-class variation via per-identity diffusion model fine-tuning [50]. While this allows for more controlled sampling, the method is computationally expensive and requires a separate SVM for each attribute. GANDiffFace, similar to SFace [4], primarily targets unconstrained images and lacks support for document-style outputs. Despite offering attribute control, GANDiffFace lacks scalability and alignment with structured document-style generation needs.
Recent studies have leveraged diffusion models for identity-consistent face generation, achieving high realism and intra-class variation. DCFace [29] combines identity and style embeddings to synthesize diverse images of the same person, while IDiff-Face [6] conditions a latent diffusion model on feature representations, achieving similar identity consistency. While effective, both methods focus on unconstrained synthesis and offer limited control over document-style formatting or demographic attributes.
The ONOT dataset [14] represents a pioneering effort in document-style face synthesis for biometric use, employing a fine-tuned Stable Diffusion [48] model and prompt engineering to generate ICAO-compliant [25] ID images. It sampled 64 images per identity across 15k identities, retaining only those with at least one compliant image, yielding usable samples for just 27% of identities. Further intra- and inter-class filtering reduced the final dataset to 255 identities. Our analysis shows that retained identities were highly similar and poorly aligned with the real identity embedding space. ONOT’s document-style control relies solely on prompts, lacking the fine-tuning needed for consistent formatting. Its demographic attribute control is rigid and not easily extensible. While ONOT shares our high-level goals, it falls short in flexibility, scalability, and diversity.
@@ 3 FLUXSynID: Document Images
Figure 1: Overview of the FLUXSynID pipeline for synthetic document image generation. Given user-defined identity attributes and constraints, descriptive prompts are generated and used to guide FLUX.1 [dev] [3] diffusion model enhanced with LoRA adapter [23] for producing realistic frontal document-style images. Please refer to Sec. 3 for details.
Frontal, standardized face images, such as those found in passports, are essential for biometric tasks such as MAD [54]. Unlike unconstrained imagery, these images must meet strict format standards: frontal pose, neutral expression, consistent lighting, and plain backgrounds. Generating such images synthetically is challenging because most diffusion or GAN-based models are trained on in-the-wild datasets. Moreover, existing methods often lack the control needed to ensure demographic diversity, leading to biased or unrealistic datasets.
Figure 1 presents an overview of the FLUXSynID pipeline for synthetic document image generation, which centers around the FLUX.1 [dev] [3] diffusion model. The process begins with user-defined identity attributes, which are used to generate natural language prompts via a large language model. These prompts guide FLUX.1 [dev], fine-tuned with a LoRA adapter [23], to produce realistic, high-resolution document-style images.
@@ 3.1 Identity Attributes
Existing face image datasets often exhibit imbalances in key identity attributes such as gender, age, and ethnicity [24]. To address this, the FLUXSynID framework allows user-defined attribute classes, each with a set of attributes with associated selection probabilities. Additionally, each class can be assigned an inclusion probability during sample generation. In Fig. 1, the gender class is always included (100%), with the Female identity attribute selected 49% of the time. Other classes, such as body type, may be included in only a portion of samples (e.g., 50%), enabling fine-grained control over attribute distributions.
To maintain semantic consistency in attribute combinations, FLUXSynID implements user-defined attribute clash detection. For example, selecting Bald as a hairstyle attribute would automatically exclude hair color attributes, as they are semantically incompatible. This system ensures the generated identity attributes remain logically coherent.
Once attributes are selected, they are stored and passed to a Large Language Model (LLM). We use Qwen2.5 [55] to generate natural language prompts that integrate all attributes in a semantically rich format. This approach aligns with FLUX.1 [dev]’s architecture, which leverages a T5 [45] text encoder to process detailed prompts effectively.
@@ 3.2 Document Image LoRA Adapter
Low-Rank Adaptation (LoRA) [23] enables efficient fine-tuning of large pre-trained models by introducing a small set of trainable parameters. Instead of updating the full model weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, the LoRA adapter keeps $W_{0}$ fixed and learns a low-rank update $\Delta W=BA$, where $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times k}$, with $r\ll\min(d,k)$. The adapted weights become $W=W_{0}+\Delta W$. This approach significantly reduces trainable parameters compared to methods such as DreamBooth [50].
We fine-tune the FLUX.1 [dev] diffusion model using a LoRA adapter to ensure stylistic consistency in generating high-resolution, vivid frontal face images with neutral expressions and a uniform white background. The adapter is trained to capture the visual style of images from the Chicago Face Database (CFD) [35, 36, 31]. Figure 2 compares results with and without LoRA, showing improvements in pose, lighting, and expression consistency.
Figure 2: (Top row): Synthetic document images generated without LoRA [23], exhibiting inconsistent head poses, facial expressions, and lighting. (Bottom row): Images generated with LoRA fine-tuning, showing more uniform, document-style outputs with neutral expressions and consistent photographic conditions.
@@ 3.3 Image Generation
Once trained, the LoRA adapter updates the weights of the pre-trained and frozen FLUX.1 [dev] model, which generates images from prompts produced by Qwen2.5. FLUX.1 [dev] is a latent diffusion model based on a rectified flow transformer [16], designed for text-to-image synthesis. It integrates Classifier-Free Guidance [22] directly into the transformer via a single modulation parameter, the guidance scale, eliminating the need for negative prompts and dual-pass inference (i.e., conditional and unconditional passes). This halves inference time and enables direct control over prompt adherence: lower guidance yields more diverse outputs, while higher guidance increases prompt adherence at the cost of variation.
As shown in Fig. 1, the generated prompt $P$ is encoded using the pre-trained and frozen CLIP-L [44] and T5-XXL [45] models. FLUX.1 [dev] combines these encoders in a complementary way. CLIP’s global embedding is fused with the timestep embedding to modulate the transformer layers. Meanwhile, T5-XXL’s token embeddings are concatenated with latent image tokens to form a unified input sequence, which is processed via self-attention. This design mimics cross-attention, enabling the text tokens to influence specific regions in the image [19].
After encoding the text prompt $P$ using both encoders ($E_{\text{CLIP-L \& T5-XXL}}$), the resulting textual conditioning $c_{P}$ is used as conditional input to FLUX.1 [dev] alongside a 16-channel Gaussian noise latent image $z_{t}$ sampled from a random seed. Through latent diffusion, FLUX.1 [dev] iteratively denoises $z_{t}$ over $t$ steps to produce the final latent image. The conditional denoising process is defined as:
@@
where DPM++ 2M [34] is used as the sampling algorithm in diffusion probabilistic modeling, and $\epsilon_{\theta^{CFD}}$ represents the FLUX.1 [dev] model with the LoRA adapter trained on CFD images (see Sec. 3.2).
The final denoised latent image, $z_{0}$, is then passed to the pre-trained FLUX.1 [dev] Variational Autoencoder (VAE) decoder network [3], $D_{\text{VAE}}$, which generates the synthetic document image, $I_{D}$, based on the textual prompt.
Figure 3: Overview of the FLUXSynID live capture image generation pipeline. Starting from a synthetic document image, three methods are used to produce identity-consistent live images: LivePortrait [20] enables controlled variations in pose and expression; Arc2Face [40] introduces natural diversity via identity embeddings; and the PuLID-based approach [21, 52] is integrated with FLUX.1 [dev] [3] via identity and edge-based conditioning, guided by modified image prompts. Please refer to Sec. 4 for details.
@@ 4 FLUXSynID: Trusted Live Capture Images
To reflect real-world identity verification scenarios (e.g., airport gates) in the FLUXSynID framework, we incorporate three complementary generative approaches: LivePortrait [20], Arc2Face [40], and PuLID [21]. As illustrated in Fig. 3, LivePortrait and Arc2Face are applied in their original forms, providing precise control and natural variability, respectively. In contrast, our primary contribution lies in the extensive adaptation of PuLID through novel conditioning strategies that enhance pose and expression diversity while preserving identity. Together, these methods generate live-style variants $I_{LL}$, $I_{LA}$, and $I_{LP}$, respectively, from the original synthetic document images $I_{D}$.
@@ 4.1 LivePortrait-based Image Generation
LivePortrait [20] is a facial animation method that leverages implicit 3D keypoints to control expression and head pose from a single image. In FLUXSynID, it is applied to each synthetic document image $I_{D}$ to generate identity-preserving live capture variants $I_{LL}$ by applying subtle, randomly sampled changes in pose and expression.
@@ 4.2 Arc2Face-based Image Generation
Arc2Face [40] generates identity-preserving facial images via ArcFace [12] identity embeddings, which condition the Stable Diffusion [48] model. The authors report superior identity similarity compared to other generative models, underscoring its effectiveness. Notably, Arc2Face requires no textual prompts, relying entirely on image-based identity embeddings. In FLUXSynID, Arc2Face is applied to the synthetic document images $I_{D}$ to produce live-style variants $I_{LA}$, complementing LivePortrait by contributing unstructured variability. This variability is driven solely by identity conditioning and helps balance LivePortrait’s more controlled outputs.